### v5

# A General Theory of Abiogenesis: Thermodynamic Selection and the Emergence of Life

Abstract
The origin of life is often approached through the lens of replication, heredity, or molecular specificity. This paper proposes a thermodynamic framework in which the emergence of life is driven by the persistence of reaction pathways that align energetically with fluctuating environmental inputs. We define a reaction viability inequality based on energy input, return, resilience, and expenditure, which selects for persistent chemical configurations without invoking heredity or genetic encoding.
We further incorporate entropic dynamics and spatial constraints into an augmented persistence function, showing that systems far from equilibrium can simultaneously increase global entropy while supporting localized chemical order. These refinements lead to the development of the Thermodynamic Abiogenesis Likelihood Model (TALM), a probabilistic extension that estimates the likelihood of persistence-driven selection under various prebiotic and planetary scenarios.
This framework redefines the conditions under which life-like organization may emerge and provides a testable, general theory for abiogenesis grounded in physical law.
1 Introduction
The origin of life remains one of the most profound and unresolved questions in science. Traditional approaches emphasize replication, heredity, and biochemical specificity under non-equilibrium thermodynamic conditions as prerequisites for natural selection. The transition from abiotic chemistry to biological evolution is often framed in terms of autocatalytic reaction networks sustained by external energy sources. However, the mechanism by which such systems become selectively retained remains underdeveloped in many existing models.
Here, we propose a general theoretical framework in which selection-like behavior can emerge in the absence of heredity or replication–driven instead by thermodynamic constraints on energy input, use, storage, and dissipation. This model formalizes a statistical pathway to abiogenesis rooted in the persistence of reaction networks whose viability depends on environmental energy availability, internal energy dynamics, and structural resilience. By reframing natural selection as a consequence of energy-compatible persistence, we unify entropy production, spatial organization, and reaction viability under a physically grounded, molecule-agnostic theory applicable across planetary environments. In this formulation, natural selection is not exclusive to genetics but can emerge in any system where energy regulation enables sustained reaction cycles.
This work contributes a scalable thermodynamic model that reconceptualizes the origin of life as the emergence of persistent chemical structure within energy-modulated environments. It requires no initial heredity, replication, or carbon-based chemistry. Grounded in first principles yet agnostic to molecular identity, the theory offers a generalizable framework for how abiogenesis may arise from prebiotic reaction chains shaped by entropy flux and diffusion constraints.
We extend this framework into a probabilistic formulation, the Thermodynamic Abiogenesis Likelihood Model (TALM), which estimates the likelihood of persistence-driven selection under diverse planetary conditions. TALM captures the emergence of selective behavior in localized chemical environments and scales across planetary and cosmic contexts. The remainder of this paper develops the model’s thermodynamic foundation, introduces its statistical structure, and explores its implications across both Earth-like and unconventional life-permitting scenarios.
As a test of the model’s predictive scope, we propose a minimal comparative system involving amphiphilic molecules of varying chain length. This system isolates entropy and structural persistence as primary variables and allows for experimental validation of TALM’s core mechanisms under well-defined aqueous conditions.
2 Terminology and Variable Definitions
To aid clarity and interdisciplinary accessibility, the following table summarizes key variables and terms used in the model:
Symbol / Term Description
$k$ The number of reactions in the reaction chain
$z(t)$ Time-varying environmental input energy available to the system. Reflects fluctuating external sources such as thermal, photonic, or chemical energy.
$y(t)$ Stored energy available within the system. Acts as a buffer during energy deficits and enhances persistence through time-lagged supply.
$x_{i}$ The energy required to perform reaction $i$.
$r_{i}$ The energy released during reaction $i$.
$s_{i}$ The energy stored in reaction products.
$\sum r_{i}$ Total energy released as usable output from individual reactions within the system. Represents internal efficiency.
$\sum x_{i}$ Total energy required to maintain all reactions in the system. Represents the system’s energetic load or metabolic cost.
$R_{n}$ Structural or network-based resilience term. Reflects robustness of the reaction topology, such as redundancy, feedback loops, or autocatalytic features.
$P_{E}$ Probability that the environment provides survivable energy fluctuations.
$P_{C}$ Probability that sufficient chemical complexity or molecular diversity exists.
$P_{P}$ Probability that reaction chains can persist over time through internal energy logic.
$P_{S}$ Probability that the system achieves dynamic stability after initial fluctuations.
$P_{life}$ Composite probability that all four preconditions are satisfied for life-like persistence.
TALM Thermodynamic Abiogenesis Likelihood Model. A generalized framework to assess the likelihood of persistent prebiotic systems under dynamic energetic constraints.
Persistence The ability of a system to sustain itself over time, despite energy fluctuations, without requiring replication or heredity.
Table 1: Summary of key variables and terms in the TALM framework.
3 Core Thermodynamic Framework
Thermodynamic Assumptions and Principles
This model is constructed on a set of thermodynamic principles and assumptions about how energy is used, transferred, and stored in prebiotic chemical systems. These assumptions are not intended to capture the full complexity of biochemistry, but rather to provide a minimalist framework capable of illustrating how selection-like dynamics may arise from energy-driven processes alone.
Energy Availability and Use
We assume that reactions in prebiotic systems are governed by basic thermodynamic constraints:
  • Every chemical reaction has an associated energy threshold (activation energy) that must be met for the reaction to proceed.
  • A system only uses as much energy as is required to complete a reaction step. If more energy is available than needed, the surplus may be either released (e.g., as heat or dissipated energy), or stored in the reaction products (e.g., as chemical potential energy).
Dynamic Environmental Input
We introduce a time-dependent energy function, $z(t)$, representing the usable energy made available from the environment. This can fluctuate due to factors such as sunlight cycles, geothermal activity, or chemical gradients. This treatment reflects empirical work showing that surface thermal cycling, particularly in hydrothermal field environments, can generate temporally structured energy availability conducive to prebiotic polymerization (Deamer, 2011) (Ross D & Deamer D, 2019). The persistence of a reaction chain is therefore not only a function of its internal energy efficiency but also its ability to maintain viability under changing energy input.
Reaction Chains as Systems
Rather than modeling isolated reactions, we conceptualize reaction chains, sequences of reactions where the output or byproducts of one reaction may fuel or enable subsequent ones. These chains can be:
  • Autocatalytic (self-amplifying)
  • Branching or linear
  • Open to energy inflow (from the environment via $z(t)$)
A chain will continue as long as the net energy available (initial input plus internal releases) meets the sum of energy requirements across the chain.
Stored Energy and Future Use
Energy stored in products ($s_{i}$) is not immediately usable unless another reaction releases it. Thus, stored energy acts as a potential buffer or reserve, contributing to a system’s future viability but not necessarily to its current energy budget.
System Definition
Let the system consist of a series of $n$ reactions, each requiring a defined amount of energy to proceed and potentially releasing or storing energy in the process. We define the following terms:
  • $k$: The number of reactions in the chain
  • $z(t)$: The total usable energy available to the system from the environment at time $t$.
  • $x_{i}$: The energy required to perform reaction $i$.
  • $r_{i}$: The energy released during reaction $i$.
  • $s_{i}$: The energy stored in reaction products.
  • $y(t)$: Residual energy at time $t$.
The Reaction Viability Inequality
For a system to persist at time $t$, the net energy budget must meet or exceed the energy costs required to sustain the chain of reactions:
$y(t)=z(t)+\sum_{i=1}^{k}r_{i}-\sum_{i=1}^{k}x_{i}$ (1)
Persistence is determined by:
$y(t)\geq 0\Rightarrow\text{System persists}$ (2)
If this condition is not met:
$y(t)<0\Rightarrow\text{System collapses}$ (3)
Here, $k$ is the number of reactions in the chain. This inequality acts as a filter: only those reaction chains whose internal energy efficiencies align with external energy fluctuations will persist. This creates an emergent selection pressure favoring systems with adaptive, energy-efficient structures.
4 Modulators of Persistence
Entropy, Diffusion, and Spatial Organization
While entropy is traditionally framed as a measure of disorder, in the context of open thermodynamic systems, it plays a more nuanced and generative role. Systems far from equilibrium, such as prebiotic chemical environments, can increase global entropy while locally producing order through self-organization and energy dissipation (Schneider & Sagan, 2005; Prigogine and Nicolis, 1977).
To visualize this, imagine entropy as a crowd entering a large hall. Left unconstrained, people spread to fill the room evenly. But if pathways, walls, or rails guide their movement, order emerges, not by suppressing motion, but by channeling it. Entropy operates similarly in chemical systems: it enables exploration of configurations, while environmental structure filters for persistence.
We propose that it is precisely this dynamic, where entropy production coexists with local structure formation, that enables the emergence of chemically diverse distributions capable of sustaining consistent reactions across fluctuating energy input. In this view, entropy is not merely a passive constraint but a fundamental driver of molecular complexity.
These entropic transformations, guided by energy flux, yield a rich combinatorial space of molecular configurations. Among these, only subsets that align energetically with environmental fluctuation will persist, thereby introducing a thermodynamic basis for selection without heredity.
Thus, entropy facilitates the exploration of chemical state space, while energy-based viability filters that space to favor persistent configurations. This interplay underpins the reaction viability inequality and reframes entropy as a formative precursor to natural selection itself.
Similarly, diffusion and spatial factors constrain the thermodynamic landscape. Reaction viability depends not only on energetic balance but also on spatial proximity of reactants, molecular crowding, and the presence of gradients or compartments that limit dispersion. Structured environments, such as mineral surfaces, vesicles, or pores, may facilitate energy retention and constrain entropy in ways that support persistence.
Extension of the Inequality with Entropic Factors
To preserve the simplicity and elegance of the original reaction viability inequality, we propose that entropy and spatial factors be treated as modulators of system persistence, influencing the feasibility of local order and structural coherence under fluctuating energy (Schneider & Sagan, 2005; Morowitz, 1968).
We introduce the term $\Delta S_{net}$ to represent the net entropy balance between the system and its environment:
$\Delta S_{net}=\Delta S_{env}-\Delta S_{local}$
where:
  • $\Delta S_{env}$ is the entropy exported to the surroundings (e.g., via heat, dispersal)
  • $\Delta S_{local}$ is the entropy reduction due to local self-organization or structure (Prigogine and Nicolis, 1977; Deamer, 2011)
For persistence to remain thermodynamically viable, $\Delta S_{net}$ must remain positive, meaning that localized decreases in entropy (e.g., formation of stable structures) are compensated by environmental entropy production (Schneider & Sagan, 2005). This relationship allows ordered reaction networks to form and persist without violating the second law.
Additionally, a spatial modifier $D(x,t)$ can be introduced to represent diffusion constraints, with low diffusion (e.g., due to compartmentalization or surface adherence) enhancing the retention of energy and structure (Rastogi, Kumar & Patnaik, 2022; Deamer & Dworkin, 2002).
To account for both entropy and diffusion constraints, we define an augmented viability function $y^{\prime}(t)$ as:
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t)$
where:
  • $y(t)$ is the original residual energy as defined earlier
  • $R_{n}$ is a resilience factor representing structural redundancy or robustness
  • $\Phi(t)$ is an entropic-diffusive penalty defined as:
    $\Phi(t)=f(\Delta S_{net}^{-1},D(x,t))$
The updated persistence condition becomes:
$y^{\prime}(t)\geq 0$
These considerations do not negate the core reaction viability inequality, but rather refine its application with physically meaningful constraints that can be applied when modeling realistic prebiotic environments. Persistence in prebiotic systems is not a purely scalar function of energy input, but a spatially and entropically modulated process, one that must be understood in the context of both energy gradients and structural boundary conditions.
5 Thermodynamic Abiogenesis Likelihood Model (TALM)
The Thermodynamic Abiogenesis Likelihood Model (TALM) extends the persistence framework developed in this paper by applying it to planetary-scale and compartmentalized chemical environments. TALM incorporates recent refinements that account for entropic transformation, diffusion constraints, and spatial organization as modulators of molecular persistence.
Entropy-Driven Exploration of Chemical Space
TALM builds on the view that entropy in open, far-from-equilibrium systems is not merely a measure of disorder, but a driver of diversity. Environmental energy fluxes, such as wet-dry cycles or thermal gradients, increase global entropy while enabling localized reductions, forming transient structures and reaction networks.
This entropic dynamism expands the chemical search space, allowing stochastic configurations that, under energy fluctuation, can be filtered via the reaction viability inequality. Thus, entropy facilitates exploration, while thermodynamics imposes a selective constraint.
Reaction Viability in Structured Environments
TALM integrates the augmented viability function:
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t)$ (4)
Here we introduce 2 scaling constants tailored to the environment:
  • $y(t)$ is the net energy balance of a reaction over time
  • $R_{n}$ is a resilience factor, representing structural or contextual redundancy
  • $\Phi(t)$ is the entropic-diffusive penalty, capturing spatial and entropy constraints
We introduce a further 2 modifiers or scaling constants in order to allow us to provide a method of weighting the entropic cost and diffusion loss:
  • $\alpha$ is the scaling constant for $1/\Delta S_{net}$ the thermodynamic cost of maintaining local order
  • $\beta$ is the scaling constant for $D(x,t)$ the diffusion penalty term factor
$\Phi(t)=\alpha\cdot\frac{1}{\Delta S_{net}}+\beta\cdot D(x,t)$
This function enables simulation and comparative evaluation across planetary contexts and captures the fact that in highly diffusive or disordered environments, persistence is penalized unless offset by compartmentalization or external entropy production.
Likelihood Formulation
We model the likelihood of at least one viable, persistent reaction configuration as:
$P_{\text{selective}}=1-(1-P(y^{\prime}(t)\geq 0))^{N}$ (5)
Where:
  • $P(y^{\prime}(t)\geq 0)$ is the probability of a reaction being viable after entropic and spatial correction
  • $N$ is the number of sampled reaction permutations in a given environment
Comparative Context: Earth and Icy Moons
To illustrate the applied potential of the TALM framework, we consider its implications for three planetary environments: early Earth, Europa, and Enceladus. Each of these contexts exhibits the environmental cycling, chemical diversity, and energy gradients necessary to support persistence-aligned reaction dynamics.
Early Earth featured diurnal heating, volcanic geothermal activity, and frequent wet-dry cycles in land-based hydrothermal fields (Deamer, 2011). These conditions likely produced spatial confinement (e.g., in clay layers, mineral pores), low diffusion rates, and fluctuating energy inputs, parameters which, within the TALM framework, increase the probability of $y^{\prime}(t)\geq 0$.
Europa, a Jovian moon, exhibits subsurface oceanic activity influenced by tidal flexing, potentially creating hydrothermal vent systems beneath an icy shell. While lacking sunlight and surface cycling, Europa’s sustained geothermal energy and possible mineral-laden ocean floor could support moderate $P_{\text{selective}}$ values, especially in enclosed or porous mineral environments.
Enceladus, with its plumes of organic-rich water vapor and evidence of internal heating, may also present transient microenvironments with localized gradients and structured surfaces. The TALM framework suggests that even episodic confinement and brief persistence windows could favor molecular configurations that align thermodynamically with available energy pulses.
This comparative perspective demonstrates how TALM can function not just as a theoretical model, but as a tool for ranking or assessing the abiogenesis potential of real-world planetary bodies.
By grounding the concept of reaction viability in energy, entropy, and structure, TALM offers a physically testable and computationally tractable framework for estimating the likelihood of abiogenesis under varying planetary conditions. It enables the comparison of chemical ecosystems by their propensity to sustain persistent molecular configurations, not through genetic inheritance, but through thermodynamic and spatial coherence.
Urability as a Prebiotic Filter
Just as habitability describes environments where life can be sustained, urability refers specifically to environments capable of supporting the origin of life. It characterizes the prebiotic window in which physical and chemical conditions–such as environmental cycling, availability of amphiphiles, and energetically favorable reaction pathways–enable the emergence of molecular complexity prior to heredity or replication Deamer (2021).
In the context of the TALM framework, urability serves as a useful pre–filter for determining which environments warrant thermodynamic modeling. While TALM evaluates whether specific reaction chains can persist under dynamic energy conditions, it assumes that basic prebiotic conditions are met. Urability helps identify planetary settings where those preconditions, such as fluctuating energy, compartmentalization, and chemical substrates, are plausible.
By combining TALM’s probabilistic approach with urability’s environmental screening criteria, we gain a two-tiered method for evaluating the likelihood of abiogenesis: first by assessing whether a planetary body is urable, and second by calculating the statistical probability of persistent reaction networks under its conditions.
6 Illustration of the Model
Reaction chain model
Figure 1 illustrates the logic of the reaction chain model under fluctuating environmental energy.
[…]
Figure 1: Flowchart of the reaction chain viability model.
Energy dynamics over time
Figure 2 illustrates energy dynamics over time, showing the interaction between stored energy and energy used in reactions..
Figure 2: Line chart showing energy dynamics over time.
Numerical Example: Reaction Chain Viability
To illustrate the viability inequality in practice, we consider a hypothetical reaction chain consisting of three sequential reactions, each with defined energy requirements ($x_{i}$), energy releases ($r_{i}$), and stored energy ($s_{i}$). We evaluate the chain’s persistence under two different environmental energy conditions, represented by $z(t)$.
Reaction Chain Parameters
We define the following reactions:
Reaction $x_{i}$ (Required) $r_{i}$ (Released) $s_{i}$ (Stored)
$R_{1}$ 8 2 1
$R_{2}$ 5 3 2
$R_{3}$ 7 1 1
The system attempts to perform all three reactions sequentially, starting with energy $z(t)$ from the environment.
Case 1: Sufficient Environmental Energy ($z(t)=10$)
  • $R_{1}$: $z=10\rightarrow z-x_{1}+r_{1}=10-8+2=4\rightarrow$ Store $s_{1}=1$
  • $R_{2}$: $z=4\rightarrow z-x_{2}+r_{2}=4-5+3=2\rightarrow$ Store $s_{2}=2$
  • $R_{3}$: $z=2\rightarrow z-x_{3}+r_{3}=2-7+1=-4\rightarrow$ Insufficient energy
Result: System collapses at step 3.
Case 2: Higher Energy Input ($z(t)=15$)
  • $R_{1}$: $z=15\rightarrow 15-8+2=9\rightarrow$ Store $s_{1}=1$
  • $R_{2}$: $z=9\rightarrow 9-5+3=7\rightarrow$ Store $s_{2}=2$
  • $R_{3}$: $z=7\rightarrow 7-7+1=1\rightarrow$ Store $s_{3}=1$
Result: Chain completes successfully. System persists and stores cumulative energy for future use.
This simplified example demonstrates the core logic of the model: system viability is governed not just by initial energy input but also by how reactions release or store energy internally. Even efficient systems may collapse under insufficient environmental support. Conversely, well-aligned chains survive and accumulate structural energy, potentially leading to selection pressure in energy-variable environments.
Illustrative Example: Thermal Cycling on Early Earth
To illustrate the applicability of the TALM framework under real environmental conditions, consider empirical measurements by Deamer and Damer (2020), who recorded daily temperature fluctuations ranging from 30–90°C in surface hydrothermal pools at Kamchatka. These fluctuations correspond to molecular-scale energy variations on the order of $10^{-20}$ to $10^{-18}$ joules per molecule. Such settings provide ideal testbeds for modeling $z(t)$, the time-dependent environmental energy input. This experimental observation is consistent with the results of Deamer and colleagues, who demonstrated that dehydration–rehydration cycles in hydrothermal fields drive non-enzymatic polymerization of nucleotide-like molecules under prebiotic conditions (DeGuzman et al, 2019) (Ross D & Deamer D, 2019).
We define $z(t)$ as a sinusoidal function representing periodic thermal cycling:
$z(t)=z_{\max}\cdot\sin(\omega t)+z_{\text{offset}}$
Where: - $z_{\max}$ is the peak energy input during daytime, - $\omega$ is the angular frequency corresponding to a full day-night cycle, - $z_{\text{offset}}$ accounts for geothermal heat or atmospheric energy retention during cooler periods.
During daytime, energy peaks may support energetically demanding reactions ($\sum x_{i}$ is high), while nighttime introduces an energy deficit. Only systems with sufficient internal storage $y(t)$, or inherently low energy requirements $x_{i}$, will remain viable during low-input phases.
For example, let a reaction system require $x_{i}=5$ energy units to maintain structural viability and release $r_{i}=1$. If $z(t)$ peaks at 6 and falls to 1 at night, then persistence through a full cycle is only possible for systems that can retain at least $y(t)\geq 3$ units of stored energy from the high-input phase.
This scenario demonstrates TALM’s central principle: differential energy storage enables selective retention. Systems capable of buffering energy during abundance become preferentially stable across environmental fluctuations. Thermal cycling, when combined with compartmentalization and diffusion-limiting substrates, thus constitutes a plausible natural engine for prebiotic selection.
Figure 3: Simulated thermal cycling and entropy export in early Earth conditions. The solid curve represents energy input $z(t)$ over a 48-hour period, modeled as a sinusoidal day–night cycle with baseline energy. The dashed curve shows entropy exported to the environment, $\Delta S_{env}(t)$, phase-shifted to reflect thermal dissipation lag. The shaded region indicates the viable persistence zone, where energy input exceeds the threshold required for reaction continuity. This illustrates how persistent systems must not only align with fluctuating $z(t)$, but also export sufficient entropy to remain thermodynamically viable.
This simplified scenario illustrates how differential energy storage enables selection: systems that store surplus energy are more likely to persist across fluctuating $z(t)$, favoring energy-aligned stability.
Furthermore, such persistence would only be thermodynamically viable if local decreases in entropy, associated with maintaining molecular structure across cycles, are offset by corresponding entropy export to the environment. In this way, day–night thermal cycling could support transient self-organization that satisfies the second law, especially in compartmentalized or surface-adhered systems where diffusion is constrained.
Applying such a model could be used to provide a new approach to the prediction of where life could originate and could be adapted to apply to hydrothermal vents, Europa’s subsurface ocean, or Titan’s hydrocarbon-rich atmosphere, or anywhere where time-variable or localized energy gradients are abundant.
Illustrative Example: Other Planetary Scenarios
In addition to early Earth, and to illustrate the flexibility of the TALM framework, we evaluate five different planetary environments. Each offers different combinations of energy cycling, chemical complexity, and structural confinement whereby we estimate the relative likelihood of life-like persistence using approximate values for energy fluctuation ($P_{E}$), chemical diversity ($P_{C}$), persistence potential ($P_{P}$), and stabilization probability ($P_{S}$). These values are illustrative and intended to demonstrate the framework’s versatility.
1. Rocky Planet with Surface Water (e.g., Early Earth)
Justification: Abundant energy sources, water as solvent, and complex chemistry.
  • $P_{E}\approx 0.9$, $P_{C}\approx 0.8$, $P_{P}\approx 0.3$, $P_{S}\approx 0.7$
$P_{life}\approx 0.1512$
2. Gas Giant Atmosphere (e.g., Jupiter)
Justification: High chemical complexity but turbulent, unstructured environment.
  • $P_{E}\approx 0.95$, $P_{C}\approx 0.7$, $P_{P}\approx 0.05$, $P_{S}\approx 0.4$
$P_{life}\approx 0.0133$
3. Icy Moon with Subsurface Ocean (e.g., Europa)
Justification: Protected liquid ocean, chemical gradients, tidal heating.
  • $P_{E}\approx 0.7$, $P_{C}\approx 0.75$, $P_{P}\approx 0.2$, $P_{S}\approx 0.6$
$P_{life}\approx 0.063$
4. Tidally Locked Exoplanet (e.g., Proxima b-like)
Justification: Persistent light/dark zones may induce stable energy differentials.
  • $P_{E}\approx 0.6$, $P_{C}\approx 0.6$, $P_{P}\approx 0.1$, $P_{S}\approx 0.5$
$P_{life}\approx 0.018$
5. Volcanically Active Super-Earth (e.g., Hypothetical Kepler Candidate)
Justification: High internal energy flux, thick atmosphere, and geochemical cycling.
  • $P_{E}\approx 0.85$, $P_{C}\approx 0.9$, $P_{P}\approx 0.25$, $P_{S}\approx 0.65$
$P_{life}\approx 0.124$
Comparative Summary
These cases demonstrate that while Earth-like planets yield high persistence probabilities, other bodies, including moons and exoplanets, fall within viable ranges, especially when energy stability and chemical diversity align. This reinforces that persistence-driven prebiotic selection is not limited to Earth-like conditions, but may emerge across a spectrum of planetary contexts where structural and thermodynamic constraints are met.
To further highlight the generalizability of the TALM framework, we summarize the core features of these environments in Table 2.
Table 2: TALM Viability Across Planetary Environments
Environment Energy Cycling Features Diffusion Constraint Estimated TALM Viability
Early Earth Wet-dry, geothermal, diurnal Mineral pores, clays Moderate to low High
Europa Tidal heating, hydrothermal Sub-ice mineral surfaces Moderate Moderate
Enceladus Geothermal, episodic plumes Porous rocky interiors Low (in confined settings) Moderate
Titan Seasonal, cryovolcanism Aerosol layers, subsurface ocean Moderate Low to moderate
Mars Geothermal (ancient), brine cycles Subsurface ice/salts Variable Low (present), moderate (past)
Venusian Clouds Photochemical, convection Aerosol droplets High Low but not negligible
Exoplanets / Exomoons Varied (e.g., tidal, rotational, orbital) Minerals, ice fractures, atmosphere Varied Case-dependent
7 Building a Simulation
Parameter Ranges and Boundary Conditions
To apply TALM in realistic simulations, we constrain the variables used in the augmented persistence inequality:
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t),\quad\text{where}\quad\Phi(t)=f(\Delta S_{net}^{-1},D(x,t))$
Based on prebiotic and physical estimates, the following parameter ranges are proposed:
  • Energy Input $E_{\text{in}}$: $10^{-21}$ to $10^{-18}$ J per molecule per event, from UV, geothermal, or chemical sources (Schrödinger, 1944; Morowitz, 1968).
  • Energy Expenditure $E_{\text{cost}}$: $10^{-20}$ to $10^{-18}$ J to sustain reaction or structural integrity.
  • Entropy Balance $\Delta S_{net}$: Net positive when $\Delta S_{env}\sim 10^{-21}$ to $10^{-20}$ J/K exceeds local ordering $\Delta S_{local}\sim-10^{-22}$ J/K.
  • Diffusion Modifier $D(x,t)$: Confined systems (vesicles, mineral pores) exhibit $10^{-14}$ to $10^{-12}$ m2/s; bulk water approximates $10^{-9}$ m2/s.
  • Resilience Term $R_{n}$: A normalized structural redundancy factor from 0 to 1; autocatalytic or feedback-rich systems may exceed 0.5.
  • Temporal Window $t$: Simulations should explore intervals from seconds to several days (e.g., 10 to $10^{5}$ seconds), reflecting real prebiotic cycling conditions.
Boundary conditions may include:
  • Fixed or variable environmental energy sources
  • Randomized initial reactant sets or network topologies
  • Compartmental geometry or surface structure modifiers
Simulation outcomes can be interpreted in terms of the probability $P(y^{\prime}(t)\geq 0)$, and aggregated across permutations $N$ to derive $P_{\text{selective}}$ for entire planetary scenarios.
Toward a Computational Model
The TALM framework lends itself naturally to simulation. A computational model can be constructed to evaluate the persistence of reaction chains under fluctuating environmental energy input $z(t)$, internal energy transformations, and thermodynamic constraints.
At each time step, the simulation proceeds as follows:
  1. 1.
    Update $z(t)$, the environmental energy input, based on a predefined or empirical function (e.g., sinusoidal, stochastic, or data-driven).
  2. 2.
    For each reaction chain:
    • Attempt reactions sequentially.
    • If the energy requirement $x_{i}$ is met, subtract it from the available energy, add the released energy $r_{i}$, and store surplus energy $s_{i}$.
    • If the energy requirement is not met, the chain halts or collapses.
  3. 3.
    Chains that persist may optionally replicate, mutate, or recombine.
  4. 4.
    The process repeats for the defined simulation window.
Future iterations may integrate entropy export ($\Delta S$), diffusion penalties, and compartmentalization effects directly into chain viability evaluations.
Baseline TALM Simulation (No Entropic or Spatial Constraints)
The following pseudocode outlines the core logic of the Thermodynamic Abiogenesis Likelihood Model (TALM) in its original, simplified form. This version models only the energy balance required to sustain a chain of prebiotic chemical reactions under a time-varying energy input $z(t)$, without accounting for entropy or diffusion penalties. Persistence is evaluated using the inequality $y(t)\geq 0$, where $y(t)$ is the system’s net residual energy after reaction execution.
1Initialize environmental energy function z(t)
2Generate a set of reaction chains
3
4For each timestep t:
5 Update environmental energy z = z(t)
6
7 For each reaction chain:
8 energy_available = z
9 Mark chain as viable
10
11 For each reaction in chain:
12 If energy_available >= required_energy:
13 Subtract required_energy from energy_available
14 Add released_energy to energy_available
15 Store any excess (optional)
16 Else:
17 Mark chain as non-viable
18 Break loop
19
20 If energy_available >= 0:
21 System persists
22 Optionally replicate or mutate
23 Else:
24 System collapses
Listing 1: Baseline TALM Simulation (No Entropy or Diffusion)
Augmented TALM Simulation with Entropy and Diffusion
This extended version of the TALM simulation incorporates entropy dynamics and diffusion-based spatial penalties into the persistence function. The augmented viability condition is given by:
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t),\quad\text{where}\quad\Phi(t)=\alpha\cdot\frac{1}{\Delta S_{\text{net}}}+\beta\cdot D(x,t)$
Entropy export ($\Delta S_{\text{env}}$) and local structural order ($\Delta S_{\text{local}}$) contribute to a net entropy balance, while diffusion $D(x,t)$ reflects the dispersal potential of the system. This model better represents physical constraints in structured environments such as pores, vesicles, or hydrothermal interfaces.
1Set constants: α (entropy scale), β (diffusion scale), γ (structure factor), T (temperature)
2
3Initialize z(t)
4Generate reaction chains with:
5 - diffusion_coefficient D(x, t)
6 - resilience R_n
7
8For each timestep t:
9 Update z = z(t)
10
11 For each reaction chain:
12 energy_available = z
13 structure_index = 0
14 Mark chain as viable
15
16 For each reaction in chain:
17 If energy_available >= required_energy:
18 Subtract required_energy from energy_available
19 Add released_energy to energy_available
20 Store stored_energy (optional)
21 Increment structure_index
22 Else:
23 Mark chain as non-viable
24 Break loop
25
26 Compute entropy terms:
27 delta_S_env = exported_heat / T
28 delta_S_local = γ * structure_index
29 delta_S_net = max(delta_S_env - delta_S_local, ε)
30
31 Compute penalty:
32 phi = α / delta_S_net + β * diffusion_coefficient
33
34 Compute augmented viability:
35 y_prime = energy_available + R_n - phi
36
37 If y_prime >= 0:
38 System persists
39 Optionally replicate or mutate
40 Else:
41 System collapses
Listing 2: Augmented TALM Simulation with Entropy and Diffusion
8 Theoretical Context and Limitations
The framework presented in this paper aligns with and extends a body of interdisciplinary work exploring the thermodynamic and information-theoretic basis for the origin of life. While traditional models of evolution emphasize replication and heredity, our approach emphasizes the prior condition necessary for such processes: the persistence of reaction systems under energetic constraints.
Limitations of the Model
While the model presented here offers a novel thermodynamic lens through which to understand the emergence of natural selection, several limitations must be acknowledged. These do not detract from the value of the conceptual framework but highlight areas for refinement, expansion, and empirical exploration.
Simplification of Chemical Complexity
This model abstracts chemical reactions into energetically discrete events, using parameters such as $x_{i}$, $r_{i}$, and $s_{i}$ to represent energy thresholds, releases, and storage. In reality, reactions involve complex molecular dynamics, intermediate states, and environmental feedbacks. No specific chemistry is modeled, and thus, while the logic is general, its direct applicability to known prebiotic chemistries remains to be validated.
Environmental Energy Model is Limited
The energy input function $z(t)$ is treated abstractly and assumed to vary arbitrarily over time. While this reflects prebiotic environments (e.g., thermal cycling, day-night radiation), no specific environmental model is encoded. Incorporating realistic environmental data or simulating z(t) from known planetary conditions could help test the plausibility of different prebiotic scenarios.
Related Works and Theories
Autocatalytic Sets and Organizational Closure
Stuart Kauffman’s theory of autocatalytic sets (Kauffman, 1993) laid the foundation for viewing life as a network of reactions that sustain and propagate themselves. His work framed emergence not as a consequence of genetic control but as a product of combinatorial chemistry and mutual reinforcement. Our model complements this by suggesting that not all autocatalytic chains persist equally, only those aligned with dynamic energy availability survive.
Dissipative Structures and Far-From-Equilibrium Systems
Ilya Prigogine’s concept of dissipative structures (Prigogine and Nicolis, 1977) posits that ordered systems can arise spontaneously in non-equilibrium environments through energy dissipation. Our framework draws from this principle by proposing that prebiotic systems capable of balancing energetic inflow and reaction requirements naturally self-stabilize, forming precursors to selection mechanisms.
Wet–dry cycles cause nucleic acid monomers to polymerize into long chains
Song et al.’s work on nucleotide polymerization (Song et al, 2019) investigates how nucleic acid monomers, when subjected to repeated wet–dry cycles under conditions simulating volcanic hot springs, can spontaneously polymerize into long chains. This study demonstrates that RNA and DNA–like oligomers can form in the absence of enzymes, highlighting a plausible non–enzymatic pathway for early genetic polymers. Their findings support the idea that environmental cycling, particularly dehydration–rehydration dynamics, can drive condensation reactions that would otherwise be thermodynamically unfavorable in continuously aqueous conditions.
Generation of oligonucleotides under hydrothermal conditions by non-enzymatic polymerization
Deguzman et al’s work on hydropthermal conditions (DeGuzman et al, 2019) studies prebiotic synthesis under fluctuating hydration that support the idea of nonequilibrium thermodynamic cycling can create and stabilize chemical complexity, without the need for enzymes or genetic templates.
Dissipative Adaptation
Jeremy England’s work on dissipative adaptation (England, 2013) formalized how driven systems statistically evolve to dissipate energy more efficiently. We align with this view, proposing that selection arises not from genetic memory, but from energy-aligned structural persistence, chains that survive fluctuations in $z(t)$ naturally dominate the prebiotic landscape.
Dynamic Kinetic Stability
Addy Pross’s theory of dynamic kinetic stability (Pross, 2012) reframes life’s origin as a transition from thermodynamic to kinetic dominance, that is, from stable molecules to stable replicative processes. Our model precedes this transition, exploring the energetic logic by which certain reaction chains become stable enough to eventually support replication.
The Hydrophobic Effect and Amphiphilic Self-Assembly
The hydrophobic effect explains the spontaneous organization of amphiphilic molecules in aqueous environments. First formalized in detail by Charles Tanford (Tanford, 1980), the effect arises from water’s tendency to exclude non-polar groups in order to preserve its hydrogen-bonding network. When hydrophobic molecules or molecular segments are introduced into water, they disrupt the local structure of the solvent, resulting in an entrop…
This behavior is central to many biological structures, including lipid membranes, and forms the physical basis for compartmentalization in the absence of genetic instruction.
Novelty of the Present Work
While these foundational theories address persistence, emergence, and order, they rarely formalize persistence in terms of a reaction viability inequality under dynamically shifting energy constraints. This work contributes a simplified but generalizable framework in which selection arises not from heredity, but from energy dynamics and reaction compatibility. It provides a thermodynamically grounded filter, one that operates even before life, as typically defined, begins.
9 Model Flexibility and Testing
While the model presented here maintains a simplified structure for clarity, it is intentionally designed to accommodate future refinements involving more complex thermodynamic behaviors.
First, the energy input function $z(t)$ may reflect environmental energy derived from various sources (e.g., photonic, thermal, or chemical). For example, a temperature-dependent function could be formulated and extended, while reaction requirements ($x_{i}$) and releases ($r_{i}$) may likewise be expressed as functions of temperature, allowing for incorporation of Arrhenius-style activation rates without disrupting the core logic of the model.
Second, stored energy is not intended to represent thermal capacity in the strict sense, but rather structural or chemical potential energy retained within the system. This distinction allows stored energy to serve as a functional buffer against environmental fluctuations, without being tied to classical specific heat or enthalpy dynamics.
Third, the model is robust under stochastic energy input. For example, if z(t) follows a Gaussian or similarly noisy distribution, stored energy acts as a compensatory term, enabling reaction chains to persist through periods of low environmental input. This introduces a selection pressure favoring systems that can store energy ahead of scarcity, a primitive resilience mechanism that parallels biological buffering strategies.
Finally, refinements to the model explicitly incorporate entropy and spatial constraints. Local entropy reduction ($\Delta S_{local}$), environmental entropy export ($\Delta S_{env}$), and diffusion penalties ($D(x,t)$) can be used to extend the core viability inequality. These modifiers reflect thermodynamic constraints in open systems and allow for more realistic modeling of self-assembly, dissipation, and compartmentalization. The inequality thus remains general but extensible, supporting both conceptual exploration and numerical simulation.
These qualities make the framework extensible. While simple in formulation, it is meant as a conceptual scaffold capable of supporting more detailed physical and chemical modeling in future work.
Test case scenario: Entropy-Driven Self-Assembly in Amphiphilic Systems
A tractable and highly illustrative system for testing TALM involves the self-assembly behavior of amphiphilic molecules in water, where structure formation is driven primarily by the hydrophobic effect. This effect arises because water molecules form hydrogen-bonded networks. When non-polar (hydrophobic) molecules are introduced, they disrupt these networks, lowering the system’s entropy. Water responds by ordering tightly around the intruding molecules. However, when many hydrophobic molecules are present, they tend to cluster together, freeing the surrounding water to return to a disordered state, thus increasing entropy.
To explore this within the TALM framework, consider a series of aqueous systems each containing 18 mL of water (1 mole) and 0.01 mole of one of three solutes: sodium acetate, sodium octanoate, and sodium dodecanoate. These molecules differ only in the length of their hydrophobic tails. When air is gently introduced, observable differences arise. Sodium acetate, which lacks a hydrophobic tail, dissolves completely with no visible change. Sodium octanoate, with a moderate tail length, forms a transient bubble. Sodium dodecanoate, with a longer hydrophobic chain, forms a stable bubble and, under microscopic inspection, many vesicles.
These behaviors are predicted by TALM as a function of chain length and the system’s ability to manage entropy and energy flow. The longer tails drive stronger self-association, lowering diffusion ($D(x,t)$) and increasing structural resilience ($R_{n}$), while also allowing water to release more entropy as structured cages collapse. Thus, the entropy cost of forming vesicles is more than offset by entropy export to the environment:
$\Delta S_{net}=\Delta S_{env}-\Delta S_{local}>0$
leading to a reduced $\Phi(t)$ and positive $y^{\prime}(t)$.
TALM predicts that structural persistence-bubbles, micelles, and vesicles-emerges at a critical chain length $L_{c}$. Based on this system:
  • For $L<C8$, self-assembly is minimal; no structures persist.
  • At $L\approx C10$, vesicle formation becomes likely.
  • Between $C10$ and $C14$, vesicles are increasingly stable.
  • Beyond $C14$, solubility may decrease, limiting structure formation.
This example shows how TALM can be experimentally tested using simple compounds and observable phenomena. It connects entropy regulation, diffusion control, and structural stability into a single framework and reinforces TALM’s core claim: persistence arises where entropy export aligns with structural coherence under fluctuating energy.
10 Theoretical Outlook
The integration of entropy production, diffusion constraints, and spatial structure into the persistence model has clarified how thermodynamic forces can shape the emergence of chemical order in prebiotic contexts. By reframing entropy not as an antagonist to organization, but as a driver of exploration within chemical state space, we offer a new lens through which the formation of persistent molecular configurations can be understood.
The TALM framework provides a means of translating this conceptual model into a probabilistic structure applicable across diverse environmental conditions. The reaction viability inequality, extended via $y^{\prime}(t)$, allows for the estimation of persistence under spatial and energetic modulation. This yields not only a theoretical basis for selection-like behavior in prebiotic chemistry but also a potential comparative lens across planetary systems.
Continued refinement of this approach, through simulation of reaction networks, laboratory analogs of compartmentalized cycling systems, and the thermodynamic profiling of planetary bodies, may illuminate the constraints and pathways most conducive to the emergence of persistent reaction architectures. In this light, the TALM framework serves not only as a theoretical advance but as a scaffold for the systematic comparison of environments where the thermodynamic roots of natural selection may have taken hold.
Moreover, pairing the TALM framework with Urability, the environmental pre-screening framework (Deamer, 2021), which assesses whether a planetary setting possesses the thermodynamic and chemical conditions necessary for prebiotic complexity, provides a two-tiered approach by first evaluating urability, then applying TALM. This provides a robust, scalable method for identifying environments most conducive to abiogenesis.
Ultimately, this model invites us to reconsider life’s origin not as a singular event but as a recurring possibility, a statistical consequence of energy flow, structure, and entropy at work within the laws of physics.
Universal Scope and Molecular Agnosticism
The Thermodynamic Abiogenesis Likelihood Model (TALM) is not confined to terrestrial assumptions of biology. It does not depend on carbon-based molecular structures, nucleotides, or biochemical replication. Rather, it defines the emergence of life-like systems in terms of energy viability and persistence within structured, entropy-driven environments. This positions TALM as a framework that is fundamentally molecularly agnostic and scalable across physical regimes.
By grounding the emergence of selection-like behavior in thermodynamic alignment rather than chemical specificity, TALM invites the consideration of exotic life forms, composed of silicon, metal-surface catalysis, quantum structures, or unknown substrates, as valid targets for origin-of-life investigation. The only constraint imposed is the existence of fluctuating energy, compartmentalizing structure, and the potential for local entropy reduction consistent with overall entropy production. This approach is consistent with the theoretical perspectives of Davies and Cleland, who argue that biology as currently known may only reflect a narrow subset of possible life-supporting structures (Cleland and Chyba, 2002; Davies, 2009).
Moreover, TALM is inherently scalable. The probability function
$P_{\text{selective}}=1-(1-P(y^{\prime}(t)\geq 0))^{N}$
allows for extrapolation from localized chemical systems ($N$ as individual reaction permutations) to whole-planetary or even galactic estimations ($N$ as the total number of habitable micro-environments or reaction domains). This scaling supports the view that abiogenesis may be not merely possible but statistically probable across large enough thermodynamic landscapes. Similar views have been proposed in the context of cosmological selection (Lineweaver et al., 2002; Walker et al., 2016), where life is treated as a common feature of entropy-utilizing complexity.
In this light, TALM does not model the origin of “life” in a strict biological sense, but rather the origin of persistent structure, thermodynamic systems that exhibit selective retention within entropic environments. It therefore provides a potential bridge between prebiotic chemistry, astrobiology, and universal principles of organized complexity.
Implications for Abiogenesis
In this framework, natural selection does not depend on heredity or replication at the outset, but emerges as a thermodynamic sieve: systems that maintain reaction viability under dynamic energy input $z(t)$, often aided by energy storage $y(t)$, network resilience $R_{n}$, and entropic compatibility, are retained, while others collapse.
This selection pressure favors not only systems that align energetically, but also those whose structure arises through entropy-driven complexity. Open systems naturally explore a vast space of molecular configurations, and entropy - through self-assembly and spontaneous ordering - acts as a creative force in generating that diversity. The energy flow then filters this space, selectively stabilizing systems that maintain coherence across energetic fluctuation.
Over time, this process preferentially retains:
  • Reaction networks that dissipate energy effectively (England, 2013)
  • Autocatalytic or structurally recursive sets (Kauffman, 1993)
  • Systems capable of storing energy or buffering against loss (Pross, 2012)
  • Structures arising from entropy-compatible self-assembly (Prigogine and Nicolis, 1977; Schneider & Sagan, 2005)
In this view, abiogenesis is not a singular event, but a gradient: a transition shaped by energy dissipation, entropic transformation, and spatial-temporal persistence. Natural selection begins not with replication, but with thermodynamic survival - an emergent property of open systems responding to energy gradients and entropic constraints.
Consequently, the TALM framework, by treating life emergence as a thermodynamic selection process, leads to several key predictive implications:
  • Life-like systems are statistically inevitable in diverse environments where fluctuating energy and chemical complexity coexist (England, 2013).
  • Non-replicative prebiotic systems may be far more common than those capable of heredity (Walker and Davies, 2013).
  • Gas giants, icy moons, and subsurface oceans could all plausibly host persistence-based chemical systems, even in the absence of traditional biosignatures (Seager, 2013).
  • Life emergence may occur in multiple distinct epochs on a single body, driven by cyclical thermodynamic phases rather than one-time events (Bains, 2004).
  • Exoplanetary systems will exhibit structure favoring thermodynamically stable bodies, potentially leading to testable correlations between exoplanet class and biopotential (Seager, 2013).
These implications suggest a revision of existing metrics of habitability, extending consideration beyond carbon-based chemistry and stable surface liquid water
11 Conclusion
The predictive application of TALM to an aqueous amphiphilic system illustrates the theory’s utility beyond abstract thermodynamic reasoning. By focusing on a simple comparative setup-acetate, octanoate, and dodecanoate under identical water and energy conditions, we show that TALM can forecast the onset of vesicle formation as a function of chain length. This experiment highlights a specific thermodynamic threshold, shaped by entropy export and diffusion constraints, where spontaneous structural persistence begins to emerge. [^1] See Appendix C for additional proposed experimental systems designed to explore structure-environment compatibility under diverse thermodynamic constraints. These findings align with classical hydrophobic theory while extending it through a persistence-focused framework.
We have presented a thermodynamic model that reframes the emergence of natural selection as a consequence of energy alignment in prebiotic reaction networks. Grounded in a reaction viability inequality, the framework captures how fluctuating environmental energy can drive the persistence, or collapse, of molecular systems, independent of heredity or replication.
By formalizing this principle with defined parameters for energy requirements ($x_{i}$), releases ($r_{i}$), storage ($s_{i}$), and network resilience ($R_{n}$), we demonstrate that differential persistence can occur even in non-replicative systems. Such a filter offers a pre-Darwinian mechanism of selection, one that operates purely through energy alignment, without heredity or reproduction.
The model is contextualized within foundational work by Kauffman, Prigogine, Deavey, England, and Pross on dissipative structures, organizational closure, and self-organization and then extended to include entropy production, diffusion constraints, and spatial modulation, enriching the model’s physical plausibility and bridging it to realistic prebiotic environments.
The Thermodynamic Abiogenesis Likelihood Model (TALM) expands this core framework to a planetary scale, allowing us to estimate the probability of persistent chemical systems arising under specific energy cycling and spatial conditions. By defining parameter ranges and comparative planetary contexts, TALM opens new avenues for assessing the likelihood of abiogenesis across a spectrum of celestial bodies.
This expansion enriches the model without compromising its principle elegance, and has a workable scope in its usage in future developments through simulation, environmental specificity, and probabilistic modeling across interstellar system settings. The incorporation of entropy and compartmentalization strengthens the physical plausibility of early self-organizing systems and when combined with environmental filters such as urability, positions the framework as a modular and physically grounded strategy for assessing the likelihood of life’s emergence across diverse planetary scenarios.
Crucially, TALM is both scalable and agnostic to molecular identity. It applies not only to terrestrial carbon-based chemistry but to any structured environment where energy gradients and entropy dynamics permit the formation of persistent configurations. Its scaling from local to galactic domains suggests that the emergence of life-like systems may be a statistical outcome of thermodynamic diversity across the universe.
This model invites a reconsideration of abiogenesis not as a rare biochemical accident, but as a recurring feature of structured, energy-driven systems. Selection may begin not with genes, but with thermodynamic compatibility, an emergent signature of matter shaped by energy and constrained by entropy.
In conclusion, we propose that selection begins not with genes, but with energy. This framework offers what may be the most physically grounded resolution yet to one of science’s most enduring mysteries: how life emerges from non-life. By identifying thermodynamic compatibility, not molecular instruction, as the initial driver of persistence, this model reframes abiogenesis as a natural consequence of energy flow, entropy modulation, and structured environments.
Appendix A: Model Parameter Ranges
Parameter Description Typical Range
$x_{i}$ Energy required per reaction $10^{-21}$ to $10^{-18}$ J
$r_{i}$ Energy released per reaction $10^{-21}$ to $10^{-18}$ J
$s_{i}$ Stored energy (buffer) Up to $r_{i}$
$R_{n}$ Network resilience 0–1 (normalized)
$\Delta S_{net}$ Net entropy balance $10^{-22}$ to $10^{-20}$ J/K
$D(x,t)$ Diffusion coefficient $10^{-14}$ to $10^{-9}$ m2/s
$z(t)$ Environmental input Sinusoidal, stochastic, or empirical
Table 3: Proposed parameter ranges for TALM model components.
Appendix B: Alternate Persistence Regimes - Experimental Proposals
TALM predicts that viable, persistent structures should emerge in systems where entropy export, energy input, and diffusion constraints are properly balanced. Below are several proposed experimental regimes to test this hypothesis beyond the aqueous amphiphile scenario:
B1. Thermal Cycling in Viscous Organic Media
  • Hypothesis: Viscosity and solvent polarity affect molecular organization and persistence.
  • Environment: Ethylene glycol or glycerol with amphiphilic or prebiotic molecules, cycled between 20–80C.
  • Prediction: Micelle-like or polymerizing structures may appear and persist through multiple cycles.
B2. Non-Aqueous Solvent Systems
  • Hypothesis: Alternative solvents may support reverse micelles or other non-lipid-based order.
  • Environment: Formamide, ethanol, or water-formamide mixtures with organic acids.
  • Prediction: Stable clusters or scaffolds may form that persist despite solvent polarity changes.
B3. Wet-Dry Cycling with Polymer Precursors
  • Hypothesis: Dehydration increases entropy flux and favors polymer formation.
  • Environment: Thin film surface hydration-dehydration using amino acids or hydroxy acids.
  • Prediction: Polymer chains persist across cycles and exhibit increased resistance to hydrolysis.
B4. Freeze-Thaw Cycles in Brine Ice Matrices
  • Hypothesis: Icy matrices may stabilize fragile structures by limiting diffusion and allowing energy buildup.
  • Environment: Slow freeze-thaw cycles in saline water or formamide solutions.
  • Prediction: Weakly bound polymers or colloids may persist and restructure after thawing.
Appendix C: Test Case Scenarios for TALM Validation
Table 4: Candidate experimental scenarios to test TALM predictions in prebiotic-like systems. Each setup is designed to probe specific parameters in the Reaction Viability Inequality.
Scenario ID Description Testable Variable(s) Prediction Based on TALM Suggested Methodology
C1 Vesicle persistence under thermocycling (e.g., sodium acetate vs. dodecanoate) Chain length, entropy gradient ($\Delta S$), hydration cycle rate Longer chains with moderate amphipathic properties will persist more stably in fluctuating entropy environments Thermocycling wet-dry system with microscopy or turbidity assays
C2 Stability of co-polymers with differing energy costs per bond Bond energy ($\Delta E$), cycle time ($\tau$) Systems with intermediate energy bonds optimized for available $\Delta E/\tau$ will show maximum persistence Synthetic polymerization assays with calorimetric energy profiling
C3 Competitive selection of reaction networks in microfluidic gradient landscapes Flow-driven $\Delta S$, competing chemical pathways More entropy-efficient or energy-compatible reactions will dominate spatially Microfluidic droplet networks with selective tracing (e.g., fluorescence)
C4 Entropy-driven assembly of amphiphilic structures Lipid concentration, gradient steepness System will preferentially select configurations with lowest dissipation per persistence unit Vesicle/micelle quantification under controlled heating gradients
C5 Chain length vs. structural stability in varying environmental entropy rates Chain length ($n$), external $\Delta S$ rate There exists an optimal range where persistence probability peaks (as per $Z>0$) Incremental chain length synthesis with stability and lifespan tracking
C6 Reversible vs. irreversible reaction dynamics in cycling environments Reaction reversibility, $\tau$, entropy change Irreversible reactions with moderate $\tau$ and favorable $\Delta E$ will outcompete reversible ones in persistence likelihood Isothermal titration calorimetry + cyclic environment simulation
Acknowledgements
The author would like to express special thanks to Dr. David Deamer and Dr. Henderson ’Jim’ Cleaves for their generous guidance, thoughtful critique, and encouragement to pursue experimental grounding for this theoretical framework. Their contributions helped refine both the conceptual clarity and the empirical relevance of this work.
Further thanks are extended to Professor Lee Cronin and Dr. Addy Pross for providing critical feedback on earlier drafts of the manuscript, which helped shape the direction and clarity of the final model.
The author acknowledges the use of generative AI tools during the drafting and editing stages of this manuscript, particularly for assistance with structuring, clarity, and technical expression. All scientific reasoning, theoretical development, and interpretation remain the responsibility of the author.